equivalent mass of metal is 2.5 times that of oxygen then mimimum molecular mass of its oxide

Grade:11

1 Answers

Arun
6 years ago
Dear Siddharth
 
Equivalent mass of oxygen = 8
Equivalent mass of metal = 2.5 * 8= 20
For minimum molecular mass, the oxide must contain one mol atom of Oxygen i.e. 16 g or two oxygens.
 
Mass of two equivalents of metal = 2* 20 = 40 g
Answer molar mass of the metal oxide = 40 +16 = 56 g
 
Regards
Arun (askIITians forum expert)